OpenStudy (anonymous):

write as improper fraction 21/3

OpenStudy (anonymous):



OpenStudy (anonymous):

7/3

OpenStudy (anonymous):

can u show me how

OpenStudy (anonymous):

i can try. you have 2 and 1/3 so the base fraction that you're working with is 1/3. 3/3=1 and you have two, so 2*(3/3) =6/3 then you add on the 1/3 that you still have standing there.

OpenStudy (anonymous):

still dont understand it

OpenStudy (anonymous):

2 is a whole number, you have to turn it into a fraction. if you have a whole pizza, that's equal to 1. so if you cut up the pizza into 8 pieces, you have 8/8 pieces of pizza or 1 whole pizza. what do you have if you eat three pieces of pizza?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

8-3=5 pieces of pizza left, so you would then have 5/8 of the original pizza.

OpenStudy (anonymous):

so all you're doing is taking your whole number and dividing it up into little pieces that match the little piece you already have (1/3) and you add them all up

OpenStudy (anonymous):

1/3 is your little piece, 1+1=2, so you can look at it like 3/3+3/3+1/3 = 7/3

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

no problem hopefully you understand it, if you don't i can try to explain it differently